I don't really know if you need the problem I was given so:
Adriana has 24 dimes and nickles in a jar. The total value of the coin in the jar is $2.10. How many of each coin does she have?


My class has been working on these type of problems as well as Digit problems for two days. I don't really understand the step after you make he chart, the solving and putting the information from the problem into equations / systems. Just for my own benefit I would really appreciate it if you can draw the chart (if possible) just so I can see if I am doing all the steps correctly. I like to ask a lot of questions, so if asking questions will take up my 5 problems , then please feel free to explain as much as you can. THANK YOU!

Answer by Alan3354(69443)

Adriana has 24 dimes and nickles in a jar. The total value of the coin in the jar is $2.10. How many of each coin does she have?
-----------
n + d = 24 (number of coins)
5n + 10d = 210 (value of coins)
---------------
n + d = 24
n +2d = 42 (2nd eqn divided by 5)
---------- Subtract
-d = -18
d = 18 dimes
----------
n = 6 nickels